what is log 4 to the base 2 raise to 1/2 + log 8 to the base 2 raised to 1/3 - log 1 to the base 4?
Answer by stanbon(75887)

what is log 4 to the base 2 raise to 1/2 + log 8 to the base 2 raised to 1/3 - log 1 to the base 4?
---------------
log2(4)^(1/2) + log2(8)^(1/3) - log4(1)
----
= (1/2)log2(4) + (1/3)log2(8) - 0
----
= (1/2)*2 + (1/3)*3 - 0
----
= 1 + 1 - 0
----
= 2
===================
Cheers,
Stan H.
==============